### v2

@@ A Proofs of Propositions 1, 2, 3, and 4
First, for every $T_{k}\in T$, we prove that every $i\in T_{k}$ most prefers his assignment $\mu(i)$ among all objects in $\omega(T_{k})$. Because the agents in $T_{k}$ trade their endowments along a cycle represented above, every agent in $T_{k}$ controls all objects in $\omega(T_{k})$. Thus, if any $i\in T_{k}$ strictly prefers some object in $\omega(T_{k})$ to $\mu(i)$, then $i$ would be able to exclusion block $\mu$, which is a contradiction. So, every $i\in T_{k}$ must most prefer $\mu(i)$ among all objects in $\omega(T_{k})$.
@@
Figure 1: Suppose there are three groups $\{1,2,3\}$, $\{4,5,6\}$, and $\{7,8,9\}$. The agents in each group trade their endowments along the cycle and most prefer their assignments among the objects owned by the group. Suppose that 22 strictly prefers 44’s endowment to his assignment, 55 strictly prefers 77’s endowment to his assignment, and 99 strictly prefers 33’s endowment of his assignment. Then, $\{2,5,9\}$ can exclusion block the allocation because they control all objects in the three cycles.
After finding $T_{1}$, we can repeat the above argument to find the group $T_{2}$ in which all agents most prefer their assignments among $O\backslash\omega(T_{1})$. Inductively applying the above argument, we can find the desired order of groups $T=(T_{1},T_{2},\ldots,T_{t})$. ∎
(1) By Proposition 3.1, when the strong core is nonempty, there exists a TTS $T^{*}=(T_{1},T_{2},\ldots,T_{t^{*}})$ such that, for every allocation $\mu$ in the strong core, every $T_{k}\in T^{*}$, and every $i\in T_{k}$, $\mu(i)\in B_{i}(O\backslash\bigcupop\displaylimits_{\ell=1}^{k-1}\omega(T_{\ell}))$. By Proposition 1, $\mu$ is in the exclusion core. Thus, the exclusion core is a superset of the strong core. Theorem 1 proves that the exclusion core is a subset of the rectified exclusion core, which is a subset of the weak core. Thus, the exclusion core is a subset of the weak core.
(2) Theorem 1 proves that the rectified exclusion core coincides with the strong core when the latter is nonempty. Since the exclusion core lies between the strong core and the rectified exclusion core, it also coincides with the strong core when the latter is nonempty.
@@
(1) Since the rectified exclusion core is nonempty, and the rectified strong core is a superset of the rectified exclusion core, which is proved below, the rectified strong core is nonempty.
Define $I_{1}=\{i\in I\backslash C:\mu(i)\in\omega(C)\}$ and $I_{2}=\{i\in I\backslash C:\omega(i)\in\mu(C)\}$. The two sets may not be disjoint. It is evident that $|I_{1}|=|I_{2}|$. Let $\sigma^{\prime}$ be any allocation such that, $\forall i\in I\backslash(I_{1}\cup C)$, $\sigma^{\prime}(i)=\mu(i)$; $\forall i\in C$, $\sigma^{\prime}(i)=\sigma(i)$; and $\forall i\in I_{1}$, $\sigma^{\prime}(i)\in\omega(I_{2})$. Then, for any $j\in I\backslash C$ such that $\mu(j)\succ_{j}\sigma^{\prime}(j)$, it must be that $j\in I_{1}$ and $\mu(j)\in\omega(C)$. Because $C_{\sigma^{\prime}\sim\mu}=C_{\sigma\sim\mu}$, for every $i\in C_{\sigma^{\prime}\sim\mu}$, it holds that $\mathcal{I}_{i}(\mu(i))\subseteq\omega(C)$. So, $\mu$ is rectification exclusion blocked by $C$ via $\sigma^{\prime}$.
The rectified strong core is a subset of the weak core, because any strong blocking coalition is a rectification blocking coalition.
The above argument can be inductively applied to each remaining $T_{k}\in T^{*}$ to prove that, for each $i\in T_{k}$, $\mu(i)\in B_{i}\big(O\backslash\bigcupop\displaylimits_{\ell=1}^{k-1}\omega(T_{\ell})\big)$. ∎
In any market, let $\mu$ be the outcome of TTC in the priority-based allocation model under some priority structure $({}_{a})_{a\in\mathcal{O}}$. Let $T=(T_{1},T_{2},\ldots,T_{t})$ denote the order of cycles removed in the procedure of TTC, where each $T_{k}$ represents the set of agents involved in the corresponding cycle. If multiple cycles are removed in the same step of TTC, their relative ranking can be arbitrary in the above order. Then, it is obvious that for each $T_{k}\in T$, $\mu(T_{k})=\omega(T_{k})$, and for each $i\in T_{k}$, $\mu(i)\in B_{i}\big(O\backslash\bigcupop\displaylimits_{\ell=1}^{k-1}\omega(T_{\ell})\big)$. By Proposition 1, $\mu$ is in the exclusion core.
Let $\mu$ be any allocation in the exclusion core. By Proposition 1, there exists a partition of agents $T=(T_{1},T_{2},\ldots,T_{t})$ such that, for every $T_{k}\in T$, $\mu(T_{k})=\omega(T_{k})$, and for every $i\in T_{k}$, $\mu(i)\in B_{i}\big(O\backslash\bigcupop\displaylimits_{\ell=1}^{k-1}\omega(T_{\ell})\big)$. Moreover, the proof of the “only if” part of Proposition 1 shows that each $T_{k}\in T$ can be chosen such that the agents in $T_{k}$ trade their endowments along a cycle. Therefore, each $T_{k}$ consists of agents who own copies of distinct object types. Then, for each $a\in\mathcal{O}$, we create a priority order a such that, for any distinct $i,j\in I_{a}$, $i{}_{a}j$ if $i\in T_{k}$, $j\in T_{k^{\prime}}$, and $k<k^{\prime}$. Since for every $T_{k}\in T$ and every $i\in T_{k}$, $\mu(i)\in B_{i}\big(O\backslash\bigcupop\displaylimits_{\ell=1}^{k-1}\omega(T_{\ell})\big)$, the outcome of TTC under the created priority structure must be $\mu$. ∎
@@ B Proofs of Lemma 1 and Theorem 1
(Unblock) Since $\mu$ is Pareto efficient, it cannot be rectification exclusion blocked by the grand coalition $I$. Now, we prove that $\mu$ cannot be rectification exclusion blocked by any coalition $C\subsetneq I$.
@@
Since $\mu$ is rectification exclusion blocked by $C$ via $\sigma$, $\mu(i^{*})\in{}^{*}(C|\omega,\mu)$. Recall that $^{*}(C|\omega,\mu)=\omega(\cup_{k=0}^{\infty}C^{k}),$ where $C^{0}=C$, and $C^{k}=C^{k-1}\cup\{i\in I\backslash C^{k-1}:\mathcal{I}_{i}(\mu(i))\subseteq\omega(C^{k-1})\}$ for every $k\geq 1$. Since $i^{*}\notin C$, there exists $\ell\geq 1$ such that $i^{*}\in C^{\ell}$ but $i^{*}\notin C^{\ell-1}$. So, $\mu(i^{*})$ is the endowment of some $i^{\prime}\in C^{\ell-1}$. However, Claim 2 below implies that $i^{*}\in C^{\ell-1}$, which is a contradiction. So, $\mu$ cannot be rectification exclusion blocked by $C$, meaning that it is in the rectified exclusion core.
The proof of Claim 2 relies on Claim 1.
@@
(1) Lemma 1 implies that the rectified exclusion core is nonempty. By Proposition 3, the rectified exclusion core is a subset of the rectified strong core, which is Pareto efficient. So, it is Pareto efficient. Below, we prove that it is equivalence-closed.
@@
(2) By Proposition 3, the rectified exclusion core is a subset of the rectified strong core, which is a subset of the weak core. So, the rectified exclusion core is a subset of the weak core. Below, we prove that it is a superset of the exclusion core.
@@
(3) When the strong core is nonempty, by Proposition 2 and Proposition 3, the exclusion core and the rectified strong core both coincide with the strong core. Since the rectified exclusion core lies between the exclusion core and the rectified strong core, it must coincide with the strong core and with the exclusion core. ∎

### v3

@@ A Proofs of Lemma 1, Proposition 1, and Proposition 2
First, for every $T_{k}\in T$, we prove that every $i\in T_{k}$ most prefers his assignment $\mu(i)$ among all objects in $\omega(T_{k})$. Because the agents in $T_{k}$ trade their endowments in a cycle represented above, every agent in $T_{k}$ controls all objects in $\omega(T_{k})$. Then, if any $i\in T_{k}$ strictly prefers some object in $\omega(T_{k})$ over $\mu(i)$, then $i$ would be able to exclusion block $\mu$ by assigning himself the preferred object and assigning his assignment $\mu(i)$ to the agent who receives his preferred object under $\mu$. This is a contradiction. So, every $i\in T_{k}$ must most prefer $\mu(i)$ among all objects in $\omega(T_{k})$.
@@
Figure 1: Suppose there are three groups $\{1,2,3\}$, $\{4,5,6\}$, and $\{7,8,9\}$. The agents in each group trade their endowments along the cycle and most prefer their assignments among the objects owned by the group. Suppose that $2$ strictly prefers $4$’s endowment over his assignment, 55 strictly prefers $7$’s endowment over his assignment, and 99 strictly prefers $3$’s endowment over his assignment. Then, because $\{2,5,9\}$ controls all objects in the three cycles, the coalition can exclusion block the allocation by letting each coalition member receive their preferred object.
After finding $T_{1}$, we can repeat the above argument to find the group $T_{2}$ in which all agents most prefer their assignments among $O\backslash\omega(T_{1})$. Inductively applying the above argument, we can find the desired order of groups, $T=(T_{1},T_{2},\ldots,T_{t})$.
(“If”) Suppose that $\mu$ is an Pareto efficient allocation and there exists a partition of agents $T=(T_{1},T_{2},\ldots,T_{t})$ such that, for every $T_{k}\in T$ and every $i\in T_{k}$, $\mu(T_{k})=\omega(T_{k})$ and $\mu(i)\in B_{i}\big(O\backslash\bigcupop\displaylimits_{\ell=1}^{k-1}\omega(T_{\ell})\big)$, and for any allocation $\sigma$ such that $I_{\sigma>\mu}\neq\emptyset$, there exists $T_{k}\in T$ such that $T_{k}\cap I_{\mu>\sigma}\neq\emptyset$ and $k<k^{\prime}$ for every $T_{k^{\prime}}\in T$ such that $T_{k^{\prime}}\cap I_{\sigma>\mu}\neq\emptyset$. We prove that $\mu$ is in the exclusion core.
[…]
(1) In the proof of the “only if” result of Lemma 1, we show that if $\mu$ is in the exclusion core, then there exists a partition of agents $T=(T_{1},T_{2},\ldots,T_{t})$ such that every $T_{k}\in T$ consists of agents who exchange their endowments in a cycle, and for every $i\in T_{k}$, $\mu(i)\in B_{i}\big(O\backslash\bigcupop\displaylimits_{\ell=1}^{k-1}\omega(T_{\ell})\big)$. This implies that $\mu$ is an outcome of TTC. Shapley and Scarf (1974) have shown that all outcomes of TTC are in the weak core. So, the exclusion core is a subset of the weak core. [^12] Alternatively, Theorem 1 proves that the exclusion core is a subset of the rectified exclusion core, which is a subset of the weak core. So, the exclusion core is a subset of the weak core.
We then prove that the exclusion core is a superset of the strong core. This obviously holds when the strong core is empty. When the strong core is nonempty, let $\mu$ be any of its elements. By Proposition 2.1, there exists a TTS $T^{*}=(T_{1},T_{2},\ldots,T_{t^{*}})$ and for every $T_{k}\in T^{*}$ and every $i\in T_{k}$, $\mu(T_{k})=\omega(T_{k})$ and $\mu(i)\in B_{i}(O\backslash\bigcupop\displaylimits_{\ell=1}^{k-1}\omega(T_{\ell}))$. For any allocation $\sigma$ such that $I_{\sigma>\mu}\neq\emptyset$, let $k^{\prime}$ be the smallest $\ell$ such that $T_{\ell}\cap I_{\sigma>\mu}\neq\emptyset$. For every $i\in T_{k^{\prime}}\cap I_{\sigma>\mu}$, since $\sigma(i)\succ_{i}\mu(i)$ and $\mu(i)\in B_{i}(O\backslash\bigcupop\displaylimits_{\ell=1}^{k^{\prime}-1}\omega(T_{\ell}))$, it must be that $\sigma(i)\in\bigcupop\displaylimits_{\ell=1}^{k^{\prime}-1}\omega(T_{\ell})$. So, there must exist $j\in\bigcupop\displaylimits_{\ell=1}^{k^{\prime}-1}T_{\ell}$ such that $\sigma(j)\notin\bigcupop\displaylimits_{\ell=1}^{k^{\prime}-1}\omega(T_{\ell})$. Let $j\in T_{k}$. So $k<k^{\prime}$. Since $T^{*}$ is a TTS, all objects that $j$ views as no worse than $\mu(j)$ are owned by $\bigcupop\displaylimits_{\ell=1}^{k}T_{\ell}$. So, $\mu(j)\succ_{j}\sigma(j)$. That is, $j\in T_{k}\cap I_{\mu>\sigma}$. Then, by the “if” result in Lemma 1, $\mu$ is in the exclusion core.
@@
In every housing market with multi-unit objects, let $\mu$ be an outcome of TTC after arbitrary tie-breaking. We prove that $\mu$ is in the exclusion core.
Let $T=(T_{1},T_{2},\ldots,T_{t})$ denote the order of cycles removed in the TTC procedure producing $\mu$, where each $T_{k}\in T$ is the set of agents involved in a cycle. If multiple cycles are removed in the same step of TTC, their relative ranking can be arbitrary in the order $T$. It is obvious that, for each $T_{k}\in T$ and each $i\in T_{k}$, $\mu(T_{k})=\omega(T_{k})$ and $\mu(i)\in B_{i}\big(O\backslash\bigcupop\displaylimits_{\ell=1}^{k-1}\omega(T_{\ell})\big)$.
Consider any allocation $\sigma$ such that $I_{\sigma>\mu}\neq\emptyset$. Let $k^{\prime}$ be the smallest $\ell$ such that $T_{\ell}\cap I_{\sigma>\mu}\neq\emptyset$. Consider any $i\in T_{k^{\prime}}\cap I_{\sigma>\mu}$. Let $x$ be the object type such that $\sigma(i)\in O_{x}$. Since $\sigma(i)\succ_{i}\mu(i)$ and $\mu(i)\in B_{i}(O\backslash\bigcupop\displaylimits_{\ell=1}^{k^{\prime}-1}\omega(T_{\ell}))$, all copies of $x$ must be assigned in the TTC procedure before $i$ is involved in a cycle. Let $J$ be the set of agents who receive the copies of $x$ under $\mu$. Since all agents in $J$ are involved in cycles before $i$, $J\cap I_{\sigma>\mu}=\emptyset$. Under $\sigma$, because $i$ receives a copy of $x$, it is impossible for all agents in $J$ to receive the copies of $x$. So, there must exist an agent $j\in J$ who receives a copy of an object type different than $x$ under $\sigma$. Since agents have strict preferences over object types and $J\cap I_{\sigma>\mu}=\emptyset$, it must be that $j\in I_{\mu>\sigma}$. This implies that $\mu$ is Pareto efficient. Let $j\in T_{k}$. So, $T_{k}\cap I_{\mu>\sigma}\neq\emptyset$ and $k<k^{\prime}$. By the “if” result of Lemma 1, $\mu$ is in the exclusion core.
The proof of the “only if” result of Lemma 1 has shown that all elements of the exclusion core are TTC outcomes. So, the exclusion core equals the set of TTC outcomes. ∎
@@ B Proofs of Lemma 2 and Theorem 1
(Unblock) Since $\mu$ is Pareto efficient, it cannot be rectification exclusion blocked by $I$. Now, we prove that $\mu$ cannot be rectification exclusion blocked by any coalition $C\subsetneq I$.
@@
Since $\mu$ is rectification exclusion blocked by $C$ via $\sigma$, $\mu(i^{*})\in{}^{*}(C|\omega,\mu)$. Recall that $^{*}(C|\omega,\mu)=\omega(\cup_{k=0}^{\infty}C^{k}),$ where $C^{0}=C$, and $C^{k}=C^{k-1}\cup\{i\in I\backslash C^{k-1}:\mathcal{I}_{i}(\mu(i))\subseteq\omega(C^{k-1})\}$ for every $k\geq 1$. Since $i^{*}\notin C$, there exists $\ell\geq 1$ such that $i^{*}\in C^{\ell}$ but $i^{*}\notin C^{\ell-1}$. So, $\mu(i^{*})$ is the endowment of some $i^{\prime}\in C^{\ell-1}$. However, Claim 2 below implies that $i^{*}\in C^{\ell-1}$, which is a contradiction. So, $\mu$ cannot be rectification exclusion blocked by $C$, meaning that it is in the rectified exclusion core. The proof of Claim 2 relies on Claim 1.
@@
(1) Lemma 2 implies that the rectified exclusion core is nonempty. It is Pareto efficient because, by Theorem 2, it is a subset of the rectified strong core, which is Pareto efficient. Below, we prove that it is equivalence-closed.
@@
(2) By Theorem 2, the rectified exclusion core is a subset of the rectified strong core, which is a subset of the weak core. So, the rectified exclusion core is a subset of the weak core. Below, we prove that it is a superset of the exclusion core.
@@
(3) When the strong core is nonempty, by Theorem 2, the rectified strong core coincides with the strong core. By Proposition 1, Theorem 2, and the result (2) proved above, the exclusion core and the rectified exclusion core lie between the strong core and the rectified strong core. So, they coincide with the strong core when it is nonempty. ∎
C Proof of Theorem 2
(1) Since the rectified exclusion core is nonempty and the rectified strong core is a superset of the rectified exclusion core, which is proved in result (2), the rectified strong core is nonempty.
Define $I_{1}=\{i\in I\backslash C:\mu(i)\in\omega(C)\}$ and $I_{2}=\{i\in I\backslash C:\omega(i)\in\mu(C)\}$. The two sets may overlap. It is evident that $|I_{1}|=|I_{2}|$. Let $\sigma^{\prime}$ be any allocation such that, $\forall i\in I\backslash(I_{1}\cup C)$, $\sigma^{\prime}(i)=\mu(i)$, $\forall i\in C$, $\sigma^{\prime}(i)=\sigma(i)$, and $\forall i\in I_{1}$, $\sigma^{\prime}(i)\in\omega(I_{2})$. Then, for any $j\in I\backslash C$ such that $\mu(j)\succ_{j}\sigma^{\prime}(j)$, it must be that $j\in I_{1}$ and $\mu(j)\in\omega(C)$. Because $C_{\sigma^{\prime}\sim\mu}=C_{\sigma\sim\mu}$, for every $i\in C_{\sigma^{\prime}\sim\mu}$, it holds that $\mathcal{I}_{i}(\mu(i))\subseteq\omega(C)$. So, $\mu$ is rectification exclusion blocked by $C$ via $\sigma^{\prime}$.
The rectified strong core is a subset of the weak core, because any strong blocking is a rectification blocking.
The above argument can be inductively applied to every remaining $T_{k}\in T^{*}$ to prove that, for every $i\in T_{k}$, $\mu(i)\in B_{i}\big(O\backslash\bigcupop\displaylimits_{\ell=1}^{k-1}\omega(T_{\ell})\big)$. ∎